If the desired inhibitor concentration is 0.2mM in a total reaction volume of 2mL, what is the volume of inhibitor you need to add to the reaction?


*this was all that was provided and was a question on a quiz I got wrong.
Answer #1

Based on the data that is available in the question I think this should be the answer.

Step 1: From the concentration and the volume provided, we find the no. of moles of inhibitor that are present in the reaction vessel by using the molarity formula and keeping in mind to convert the concentration from micromolars to molar.

Step 2: We use the formula CONCENTRATION = (NO. OF MOLES) / VOLUME to find the volume of inhibitor

STEP 1:

Molarity = moles of solute(n) / volume of solution in litres

0.2mM = n / 2mL

2 x 10-7 M = n / 2 x 10-3 L

n= 4 x 10-10 moles

STEP 2:

CONCENTRATION = (NO. OF MOLES) / VOLUME(V)

2 x 10-7 M = 4 x 10-10 moles / V

V = 4 x 10-10 moles / 2 x 10-7 moles L-1 {since molarity can be expressed in moles per litre}

V = 0.002 L of Inhibitor should be added.